OpenStudy (anonymous):

A monument is built out of toothpicks. There are 25 rows and each row has eight more toothpicks than the row above. The top row has five toothpicks. How many toothpicks are on the bottom row?

OpenStudy (unklerhaukus):

top row has 5 =5 second row has (8+5) =13 third row has 8+(8+5) =18 fourth row has 8+(8+8+5) =21 ,... nth row has 8(n-1)+5